Status Report: 30 November 2005
Wednesday Afternoon ---
A BIG storm is approaching the ARB with 3-5 inches of liquid expected to fall starting this evening and lasting at least 24 hours. It will be warm at the start (0C at 7.5 kft) and cold at the end (0C at 4.5 kft).
Both scanning research radars should begin collecting data no later than 6 pm PST, and be prepared to continue for more than 24 hours.
SMART-R is up and running.
X-band is waiting for crew to arrive this afternoon.
No sounding system at Slough House until week of 12DEC.; no supplemetal sondes from OAK needed before then.
Installation of S-band profiling radars will be delayed due to engineers being stranded in Nevada yesterday.
Late Wednesday Afternoon
Kingsmill and Ayers arrived at the X-band around 1 pm MST. The radar appears to be ok, but the phone is dead.
X-band has been operational and collecting data since 01 UTC 1 December. Data is being shipped to Boulder. Work on displaying this data is in progress---stay tuned.